# Inference for dynamic Erdős-Rényi random graphs under regime switching

### v1

@@ Abstract.
This paper examines a model involving two dynamic Erdős-Rényi random graphs that evolve in parallel, where the edges in each of these graphs alternate between being present and absent according to specific on- and off-time distributions. In our setup there is regime switching, in that the graph that we observe depends on the mode of the background process at that point in time, this background process being modeled as an alternating renewal process.
In our setup we only have access to aggregate quantities such as the number of active edges or the counts of specific subgraphs (such as stars or complete graphs) in the observed graph; importantly, we do not observe the mode. The objective is to estimate the on- and off-time distributions of the edges in each of the two dynamic Erdős-Rényi random graphs, as well as the distribution of time spent in each of the two modes. By employing parametric models for the on- and off-times and the background process, we develop a method of moments approach to estimate the relevant parameters. Numerical experiments are conducted to demonstrate the effectiveness of the proposed method in recovering these parameters.
@@
Acknowledgments. This research was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement no. 945045, and by the NWO Gravitation project NETWORKS under grant agreement no. 024.002.003. [Uncaptioned image] Date: January 20, 2025.
@@ 1. Introduction
In the setup that we consider we assume that our system is in stationarity. This concretely entails that each of the edges in graph $i$ probabilistically behaves as a (static) Erdős-Rényi random graph with $N$ vertices and ‘on-probability’
@@
A concrete example of our estimation problem is: in a setting in which we know that $X_{i}$ has a geometric distribution with parameter $p_{i}$, $Y_{i}$ a geometric distribution with parameter $q_{i}$ and $Z_{1},Z_{2}$ geometric distributions with parameter $p_{0}$ and $q_{0}$, respectively. We wish to estimate these six parameters based on the evolution of the number of triangles in the observed graph.
Subgraph counts. As mentioned above, the input of our estimation procedure corresponds to a time series that recors the evolution of specific subgraph counts in the observed graph. In the sequel we restrict ourselves to the following types of subgraphs:
@@
    A star $S_{\ell}$ with $\ell\in\{2,3,\ldots\}$ is a tree with one ‘internal vertex’ and $\ell-1$ ‘leaves’. Thus $S_{2}$ is an edge and $S_{3}$ is a line or (equivalently) a wedge.
@@
$F_{i}(k):=\mathbb{P}(X_{i}\leqslant k),\qquad\qquad G_{i}(k):=\mathbb{P}(Y_{i}\leqslant k)\qquad\qquad H_{i}(k):=\mathbb{P}(Z_{i}\leqslant k)\,.$
@@
$f_{i}(k):=\mathbb{P}(X_{i}=k),\qquad\qquad g_{i}(k):=\mathbb{P}(Y_{i}=k)\qquad\qquad h_{i}(k):=\mathbb{P}(Z_{i}=k)\,.$
@@
$\bar{f}_{i}(k):=\frac{\sum_{\ell=k}^{\infty}f_{i}(k)}{\mathbb{E}\,X_{i}},\qquad\qquad\bar{g}_{i}(k):=\frac{\sum_{\ell=k}^{\infty}g_{i}(k)}{\mathbb{E}\,Y_{i}}\qquad\qquad\bar{h}_{i}(k):=\frac{\sum_{\ell=k}^{\infty}h_{i}(k)}{\mathbb{E}\,Z_{i}}\,,$
@@
    As mentioned, the observations correspond to subgraph counts. These could concern in principle any subgraph, but we explicitly work out the cases of the number of edges, complete graphs, and stars (or any combination of these). Setting up the moment equations requires the evaluation of specific moment. We derive closed-form expressions for specific types of (cross) moments, which form a crucial element in our approach.
@@
    The performance of our estimation procedure is assessed through a series of experiments. Notably, our approach can deal with relatively high numbers of parameters; in one of the examples we succeed in estimating as many as seven unknown parameters. Our experiments also shed light on the effect of the choice of the type of subgraph on the standard deviation of the estimator.
@@ 2. Literature review
Methodologies. In various studies methods have been devised for estimating the unobservable modulating process from observations of the process that it influences; see e.g. the survey [2] in the context of queues. An example is [5], in which the bull and bear market states, along with transitions between them, are modeled using a Markov switching framework, with parameters that are estimated by likelihood maximization. Another approach is followed in [4], considering a setup in which packet arrival processes are modeled as Markov-modulated processes, with the parameters being inferred by matching characteristics of the observed process, such as the mean arrival rate, the variance-to-mean ratio, and the third moment of arrival counts. Similarly, [3] examines a population process with independent sojourn times driven by a Markov-modulated Poisson process arrival mechanism. In queueing theory, this process corresponds to the M/M/$\infty$ queue in a random environment, as discussed in [11]. These models are often interpreted as birth-death processes under modulation, with applications across various domains [14, 1, 11]. For such models, an Expectation-Maximization algorithm is frequently employed to obtain maximum likelihood estimates. In communication networks, [13] develops a methodology based on a hidden Markov model (HMM) so as to infer network states (e.g., congested or non-congested) by observing packet losses and delays in communication channels. Parameter estimation for the HMM is performed using the Expectation-Maximization algorithm. Similarly, [10] utilizes a HMM trained in a supervised learning framework to analyze vibration signals from bearings, enabling health monitoring and fault prediction. In the context of gene regulation, [14] assumes Gamma-distributed on and off times for genes, estimating parameters by matching the first three moments with their empirical counterparts.
None of the frameworks described above covers our setting of dynamic random graphs under regime switching. Our approach is based on the method of moments, incorporating moments that capture transitions between subsequent snapshots. Importantly, we estimate parameters for both the hidden underlying process and the observed process.
@@ 3. Method of moments
In this section, we propose an estimator for the unknown parameters based on the method of moments. The main idea of the method of moments is that we equate (i) theoretical moment expressions (i.e., expressions for certain expectations, in terms of the parameters to be estimated), and (ii) their empirical counterparts (i.e., estimates of these expectations, as obtained from observations). This leads to a number of equations, from which estimates of the unknown parameters can be derived. Evidently, to use this methodology it is essential that the number of moment equations is sufficiently large so as to be able to estimate the unknown parameters.
@@
for $i=1,2.$ Here $\varrho_{i}$ is to be interpreted as the probability that in stationarity the edge in graph $i$ is present, and the factor $1-\bar{f}_{i}(1)$ represents the probability that it remains present. The probability that, in graph $i$, a specific edge is present at time $k$ and another edge at time $k+1$, is given by $\varrho_{i}^{2}$, with $i=1,2$, by virtue of each graph’s edges evolving independently. Since these probabilities differ, in our computations of the cross moments it is necessary to distinguish these cases.
@@ 3.1. Edges
This decomposition can be understood as follows. The first term in the right-hand side of Equation (2) corresponds to the scenario where at time $k$ and time $k+1$ two different graphs are observed. Since each graph contains (${N\choose 2}$ edges, this can occur in (${N\choose 2}\times{N\choose 2}$ possible ways. The probability of observing a specific edge at both times $k$ as well as $k+1$ is given by $\varrho_{1}\,\varrho_{2}$, as a consequence of the assumed independence and stationarity. The second term in the right-hand side of Equation (2) corresponds to the scenario where graph 1 is observed at time $k$ as well as at time $k+1$. The probability of a specific edge being present at both time instances is $\varrho_{1}(1-\bar{f}_{1}(1))$ if the edge is the same (hence the weight $a_{2}$), and otherwise this probability is $\varrho_{1}^{2}$ (with the weight $a_{1}$). This reasoning also applies to the third term in (2), covering the scenario where it is graph 2 that is observed at the time instances $k$ and $k+1$.
@@ 3.3. Stars
Figure 1. Illustration of common edges for $S_{4}$ and $\tilde{S}_{4}$ when $m=2$ (left panels) and $m=3$ (right panels), where the grey lines and dots represent ${S}_{4}$, and the blue lines and dots represent $\tilde{S}_{4}$.
@@
$a_{m}(\ell+m)(\ell-m)$
cases in which the two graphs at $k$ and $k+1$ sharing $m$ vertices have no edges in common. The probability of observing the two stars is $\varrho_{1}^{\ell-1}\,\varrho_{1}^{\ell-1}$. Again appealing to Proposition 1, there are $a_{m}\,m(m-1)$ cases corresponding to the scenario where the two graphs at $k$ and $k+1$ sharing $m$ vertices have $1$ edge in common. The probability of observing two stars in this scenario is $\varrho_{1}^{\ell-1}\,\varrho_{1}^{\ell-2}\,(1-\bar{f}_{1}(1))$. There are $a_{m}m$ cases corresponding to the scenario where two stars sharing $m$ vertices have $m-1$ edges in common. The probability of observing these two stars is $\varrho_{1}^{\ell-1}\,\varrho_{1}^{\ell-m}\,(1-\bar{f}_{1}(1))^{m-1}$. Adding up these three parts, and multiplying by $r_{1}$, results in the second term in the right-hand side of (4). The third term in Equation (4), corresponding to the scenario where graph 2 is observed at $k$ and $k+1$, is derived along the same lines.
@@ 3.4. Estimator
Now that we have expressions for the single snapshot moments and the cross moments, we can propose our estimator. Suppose for the moment that there is a single unknown parameter for each of the random variables involved, so that there is a total of six unknown parameters that we would like to estimate.
@@ 4. Numerical examples
We illustrate the performance of our estimation approach through a series of numerical experiments. Specifically, in these experiments we let $Z_{1}\sim\mathbb{G}(p_{0}),\,Z_{2}\sim\mathbb{G}(q_{0})$ with $p_{0}=0.3,\,q_{0}=0.6$, and we consider the following three instances:
@@
    Case II: $X_{1}\sim\mathbb{W}(1.5,\alpha),\,Y_{1}\sim\mathbb{G}(q_{1}),\,X_{2}\sim\mathbb{W}(1.5,\beta),\,Y_{2}\sim\mathbb{G}(q_{2})$, with the same unknown parameters as in Case I. This makes again six parameters, but now we use the moment equations based on $\mathcal{A}_{T,0}$, $\mathcal{S}_{T,3,0}$, $\mathcal{S}_{K,4,0}$ and $\mathcal{A}_{T,1}$, $\mathcal{S}_{T,3,1}$, $\mathcal{S}_{K,4,1}$;
@@
General conclusions from Figures 2, 3, and 4 are that (i) all parameters are reproduced with a relatively high level of precision, and (ii) the bell-shaped curves suggest that the estimators may be approximately Gaussian.
We continue by presenting more detailed observations, starting with a comparison between Case I and Case II. It is seen that the estimates obtained for Case II have smaller standard deviations than their counterparts for Case I. In Case I, we use triangles ($K_{3}$), while in Case II we use 4-stars (i.e., $S_{4}$). They both require three edges, but the number of triangles is, under our parameter setting, substantially lower than 4-stars; it may be this low number of triangles in Case I that makes the estimator perform less well. (For example, in a single run, the mean number of $K_{3}$ is about 75, whereas the mean number of $S_{4}$ is about 900.)

### v2

@@ Abstract.
This paper examines a model involving two dynamic Erdős–Rényi random graphs that evolve in parallel, with edges in each graph alternating between being present and absent according to specified on- and off-time distributions. A key feature of our setup is regime switching: the graph that is observed at any given moment depends on the state of an underlying background process, which is modeled as an alternating renewal process. This modeling framework captures a common situation in various real-world applications, where the observed network is influenced by a (typically unobservable) background process. Such scenarios arise, for example, in economics, communication networks, and biological systems.
In our setup we only have access to aggregate quantities such as the number of active edges or the counts of specific subgraphs (such as stars or complete graphs) in the observed graph; importantly, we do not observe the mode. The objective is to estimate the on- and off-time distributions of the edges in each of the two dynamic Erdős-Rényi random graphs, as well as the distribution of time spent in each of the two modes. By employing parametric models for the on- and off-times and the background process, we develop a method of moments approach to estimate the relevant parameters. Experimental evaluations are conducted to demonstrate the effectiveness of the proposed method in recovering these parameters.
@@
Acknowledgments. This research was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement no. 945045, and by the NWO Gravitation project NETWORKS under grant agreement no. 024.002.003. [Uncaptioned image] Date: May 26, 2025.
@@ 1. Introduction
In the setup that we consider, we assume that our system is in stationarity. This concretely entails that each of the edges in graph $i$ probabilistically behaves as a (static) Erdős-Rényi random graph with $N$ vertices and ‘on-probability’
@@
A concrete example of our estimation problem is: in a setting in which we know that $X_{i}$ has a geometric distribution with parameter $p_{i}$, $Y_{i}$ a geometric distribution with parameter $q_{i}$ and $Z_{1},Z_{2}$ geometric distributions with parameters $p_{0}$ and $q_{0}$, respectively. We wish to estimate these six parameters based on the evolution of the number of triangles in the observed graph.
Subgraph counts. As mentioned above, the input of our estimation procedure corresponds to a time series that records the evolution of specific subgraph counts in the observed graph. In the sequel we restrict ourselves to the following types of subgraphs:
@@
    A star $S_{\ell}$ with $\ell\in\{2,3,\ldots\}$ is a tree with one ‘internal vertex’ and $\ell-1$ ‘leaves’. Thus $S_{2}$ is an edge $S_{2}$ is an edge, and $S_{3}$ is a path of length two, also known as a wedge in an undirected graph.
@@
We focus on three fundamental structural patterns since they offer meaningful insights across a wide range of application domains.
  • Edge. An edge represents a direct connection between two nodes and serves as the basic unit of interaction in a network. In social networks, for instance, edges correspond to friendships or connections, where a greater number of edges typically indicates denser interactions, which is characteristic of tightly connected communities.
  • Complete subgraph. The presence of complete subgraphs typically signifies tightly knit groups, such as close friend circles or collaborative teams. In social and communication networks, complete subgraphs are often associated with trust and mutual engagement. Platforms like Facebook and LinkedIn leverage this to suggest groups. Conversely, fraudulent or adversarial networks often lack such closed patterns, tending instead toward more random or sparse connectivity. In infrastructure networks, triangles indicate redundant paths, contributing to greater robustness and fault tolerance.
  • Star. Stars are used to capture the potential for new connections and to identify central or influential nodes. In social networks, stars support recommendation systems such as Facebook’s ‘People You May Know’ and LinkedIn’s connection suggestions. They also help uncover key individuals who serve as bridges between different communities. In financial transaction networks, a star-like structure (such as when User X transfers money to Y, and Y to Z, without a direct link between X and Z) may signal anomalous behavior and merit further investigation.
@@
$F_{i}(k):=\mathbb{P}(X_{i}\leqslant k),\qquad\qquad G_{i}(k):=\mathbb{P}(Y_{i}\leqslant k),\qquad\qquad H_{i}(k):=\mathbb{P}(Z_{i}\leqslant k)\,.$
@@
$f_{i}(k):=\mathbb{P}(X_{i}=k),\qquad\qquad g_{i}(k):=\mathbb{P}(Y_{i}=k),\qquad\qquad h_{i}(k):=\mathbb{P}(Z_{i}=k)\,.$
@@
$\bar{f}_{i}(k):=\frac{\sum_{\ell=k}^{\infty}f_{i}(\ell)}{\mathbb{E}\,X_{i}},\qquad\qquad\bar{g}_{i}(k):=\frac{\sum_{\ell=k}^{\infty}g_{i}(\ell)}{\mathbb{E}\,Y_{i}}\qquad\qquad\bar{h}_{i}(k):=\frac{\sum_{\ell=k}^{\infty}h_{i}(\ell)}{\mathbb{E}\,Z_{i}}\,,$
@@
    As mentioned, the observations correspond to subgraph counts. These could concern in principle any subgraph, but we explicitly work out the cases of the number of edges, complete graphs, and stars (or any combination of these). Setting up the moment equations requires the evaluation of a specific moment. We derive closed-form expressions for specific types of (cross) moments, which form a crucial element in our approach.
@@
    The performance of our estimation procedure is assessed through a series of experiments. Notably, our approach can deal with relatively high numbers of parameters; in one of the examples, we succeed in estimating as many as seven unknown parameters. Our experiments also shed light on the effect of the choice of the type of subgraph on the standard deviation of the estimator.
@@ 2. Literature review
Methodologies. In various studies methods have been devised for estimating the unobservable modulating process from observations of the process that it influences; see e.g. the survey [2] in the context of queues. An example is [5], in which the bull and bear market states, along with transitions between them, are modeled using a Markov switching framework, with parameters that are estimated by likelihood maximization. Another approach is followed in [4], considering a setup in which packet arrival processes are modeled as Markov-modulated processes, with the parameters being inferred by matching characteristics of the observed process, such as the mean arrival rate, the variance-to-mean ratio, and the third moment of arrival counts. Similarly, [3] examines a population process with independent sojourn times driven by a Markov-modulated Poisson process arrival mechanism. In queueing theory, this process corresponds to the M/M/$\infty$ queue in a random environment, as discussed in [11]. These models are often interpreted as birth-death processes under modulation, with applications across various domains [14, 1, 11]. For such models, an Expectation-Maximization algorithm is frequently employed to obtain maximum likelihood estimates. In communication networks, [13] develops a methodology based on a hidden Markov model (HMM) so as to infer network states (e.g., congested or non-congested) by observing packet losses and delays in communication channels. Parameter estimation for the HMM is performed using the Expectation-Maximization algorithm. Similarly, [10] utilizes an HMM trained in a supervised learning framework to analyze vibration signals from bearings, enabling health monitoring and fault prediction. In the context of gene regulation, [14] assumes Gamma-distributed on and off times for genes, estimating parameters by matching the first three moments with their empirical counterparts.
None of the frameworks described above covers our setting of dynamic random graphs under regime switching. Our problem is distinct due to the following features:
  • We do not observe individual edge dynamics; instead, we only observe aggregate information, i.e. the total number of active edges at each time point. This aggregation renders many standard methods inapplicable.
  • The on and off-times of each edge can follow arbitrary distributions, not restricted to any particular parametric family.
To the best of our knowledge, no existing method can handle this combination of features. Our approach is based on the method of moments, incorporating moments that capture transitions between subsequent snapshots. Importantly, we estimate parameters for both the hidden underlying process and the observed process.
@@ 3. Method of moments
In this section, we propose an estimator for the unknown parameters based on the method of moments. The main idea of the method of moments is that we equate (i) theoretical moment expressions (i.e., expressions for certain expectations, in terms of the parameters to be estimated), and (ii) their empirical counterparts (i.e., estimates of these expectations, as obtained from observations). This leads to a number of equations, from which estimates of the unknown parameters can be derived. Evidently, to use this methodology, it is essential that the number of moment equations is sufficiently large so as to be able to estimate the unknown parameters.
@@
for $i=1,2.$ Here $\varrho_{i}$ is to be interpreted as the probability that in stationarity the edge in graph $i$ is present, and the factor $1-\bar{f}_{i}(1)$ represents the probability that it remains present. The probability that, in graph $i$, a specific edge is present at time $k$ and another edge at time $k+1$, is given by $\varrho_{i}^{2}$, with $i=1,2$, by virtue of each graph’s edges evolving independently. Since these probabilities differ, in our computations of the cross moments, it is necessary to distinguish these cases.
@@ 3.1. Edges
This decomposition can be understood as follows. The first term in the right-hand side of Equation (2) corresponds to the scenario where at time $k$ and time $k+1$ two different graphs are observed. Since each graph contains (${N\choose 2}$ edges, this can occur in (${N\choose 2}\times{N\choose 2}$ possible ways. The probability of observing a specific edge at both times $k$ as well as $k+1$ is given by $\varrho_{1}\,\varrho_{2}$, as a consequence of the assumed independence and stationarity. The second term in the right-hand side of Equation (2) corresponds to the scenario where graph 1 is observed at time $k$ as well as at time $k+1$. The probability of a specific edge being present at both time instances is $\varrho_{1}(1-\bar{f}_{1}(1))$ if the edge is the same (hence the weight $a_{2}^{(2)}$), and otherwise this probability is $\varrho_{1}^{2}$ (with the weight $a_{1}^{(2)}$). This reasoning also applies to the third term in (2), covering the scenario where it is graph 2 that is observed at the time instances $k$ and $k+1$.
@@ 3.3. Stars
Figure 1. Illustration of common edges for $S_{4}$ and $\tilde{S}_{4}$ when $m=2$ (left panels) and $m=3$ (right panels), where the grey edges and nodes represent ${S}_{4}$, and the blue edges and nodes represent $\tilde{S}_{4}$.
@@
$a_{m}^{(\ell)}(\ell+m)(\ell-m)$
cases in which the two graphs at $k$ and $k+1$ sharing $m$ vertices have no edges in common. The probability of observing the two stars is $\varrho_{1}^{\ell-1}\,\varrho_{1}^{\ell-1}$. Again appealing to Proposition 1, there are $a_{m}^{(\ell)}\,m(m-1)$ cases corresponding to the scenario where the two graphs at $k$ and $k+1$ sharing $m$ vertices have $1$ edge in common. The probability of observing two stars in this scenario is $\varrho_{1}^{\ell-1}\,\varrho_{1}^{\ell-2}\,(1-\bar{f}_{1}(1))$. There are $a_{m}^{(\ell)}m$ cases corresponding to the scenario where two stars sharing $m$ vertices have $m-1$ edges in common. The probability of observing these two stars is $\varrho_{1}^{\ell-1}\,\varrho_{1}^{\ell-m}\,(1-\bar{f}_{1}(1))^{m-1}$. Adding up these three parts, and multiplying by $r_{1}$, results in the second term in the right-hand side of (4). The third term in Equation (4), corresponding to the scenario where graph 2 is observed at $k$ and $k+1$, is derived along the same lines.
@@ 3.4. Estimator
Now that we have expressions for the single snapshot moments and the cross moments, we can propose our estimator. Suppose for the moment that there is a single unknown parameter for each of the random variables involved, so that there is a total of six unknown parameters that we would like to estimate. Theoretically, we require six independent moments involving the six unknown parameters, where in principle any value of $\ell$ could be used. However, smaller values of $\ell$ (a) lead to simpler expressions for the moments, and (b) autocorrelations that are more clearly separated from 0. Therefore, in our experimental evaluation, we focused on $\ell\in{2,3,4}$.
@@ 4. Numerical examples
We illustrate the performance of our estimation approach through a series of experimental evaluations. Specifically, in these experiments we let $Z_{1}\sim\mathbb{G}(p_{0}),\,Z_{2}\sim\mathbb{G}(q_{0})$ with $p_{0}=0.3,\,q_{0}=0.6$, and we consider the following three instances:
@@
    Case II: $X_{1}\sim\mathbb{W}(1.5,\alpha),\,Y_{1}\sim\mathbb{G}(q_{1}),\,X_{2}\sim\mathbb{W}(1.5,\beta),\,Y_{2}\sim\mathbb{G}(q_{2})$, with the same unknown parameters as in Case I. This makes again six parameters, but now we use the moment equations based on $\mathcal{A}_{T,0}$, $\mathcal{S}_{T,3,0}$, $\mathcal{S}_{T,4,0}$ and $\mathcal{A}_{T,1}$, $\mathcal{S}_{T,3,1}$, $\mathcal{S}_{T,4,1}$;
@@
Each run consists of a simulation phase followed by an estimation phase, both executed within the MATLAB environment. In each run, we simulate the process once, record the occurrences of $K_{2}$, $K_{3}$, $S_{3}$, and $S_{4}$, and compute both the single-snapshot moments and the cross-temporal moments. We then apply the two-step estimation procedure described in Section 3.4 to estimate the corresponding unknown parameters. The simulation phase takes approximately 44 seconds, while the estimation phase never exceeds 5 seconds.
General conclusions from Figures 2, 3, and 4 ($N=15$) are that (i) all parameters are reproduced with a relatively high level of precision, and (ii) the bell-shaped curves suggest that the estimators may be approximately Gaussian. For comparison, we also present the estimates for $N=15$ and $N=30$ in Table 2.
Statistic $N=15$ $N=30$
Case I $\widehat{p}_{0}^{(L)}$ $0.2999\,(0.0020)$ $0.3000\,(0.0017)$
$\widehat{q}_{0}^{(L)}$ $0.6002\,(0.0030)$ $0.6000\,(0.0027)$
$\widehat{\alpha}^{(L)}$ $0.5003\,(0.0127)$ $0.5014\,(0.0214)$
$\widehat{q}_{1}^{(L)}$ $0.4000\,(0.0059)$ $0.4001\,(0.0111)$
$\widehat{\beta}^{(L)}$ $0.3000\,(0.0044)$ $0.3005,\,(0.0070)$
$\widehat{q}_{2}^{(L)}$ $0.8001\,(0.0285)$ $0.8028\,(0.0444)$
Case II $\widehat{p}_{0}^{(L)}$ $0.3000\,(0.0019)$ $0.2999\,(0.0017)$
$\widehat{q}_{0}^{(L)}$ $0.6001\,(0.0028)$ $0.6000\,(0.0027)$
$\widehat{\alpha}^{(L)}$ $0.5004\,(0.0114)$ $0.5013\,(0.0203)$
$\widehat{q}_{1}^{(L)}$ $0.4000\,(0.0059)$ $0.4001\,(0.0104)$
$\widehat{\beta}^{(L)}$ $0.3001\,(0.0035)$ $0.3004\,(0.0060)$
$\widehat{q}_{2}^{(L)}$ $0.8006\,(0.0226)$ $0.8022,\,(0.0381)$
Case III $\widehat{\alpha}^{(L)}$ $0.5322\,(0.1660)$ $0.8304\,(0.5848)$
$\widehat{\lambda}^{(L)}$ $1.5232\,(0.3238)$ $1.4089\,(0.6460)$
[…]
We continue by presenting more detailed observations, starting with a comparison between Case I and Case II. It is seen that the estimates obtained for Case II have smaller standard deviations than their counterparts for Case I. In Case I, we use triangles ($K_{3}$), while in Case II, we use 4-stars (i.e., $S_{4}$). They both require three edges, but the number of triangles is, under our parameter setting, substantially lower than 4-stars; it may be this low number of triangles in Case I that makes the estimator perform less well. (For example, in a single run, the mean number of $K_{3}$ is about 75, whereas the mean number of $S_{4}$ is about 900.)